OpenStudy (anonymous):

What number is 73% of 215? (1 point) a.) 215 = 0.73n b.) n = 0.73 × 215 c.) 0.73 = n × 215 d.) 215 = 73n

OpenStudy (xapproachesinfinity):

what number is translated as unknown let it be x "is" is translated as equal so x= 73%of 215 is translated as 73/100 times 215 so \(\large x=\frac{73}{100}\times215\)

OpenStudy (xapproachesinfinity):

Do the rest to get what you want

OpenStudy (anonymous):

73% of 215=73/100x215=0.73x215